Hi, Can you please help with this equation? I have a test first thing tomorrow morning and I am very stuck. I know what the answers should be, (x=1.5 and y=-0.8) but and having trouble getting there.
The equation is as follows:
1.2(x+4)-2.5(3y-2) = 2
4/5(3x-5)+ 3/2(3y-4) = -10
This is what I've got so far:
1.2(x+4)-2.5(3y-2) = 2
.08(3x-5)+1.5(3y-4) = -10
1.2x + 4.8 -7.5y+5 = 2
2.4x-4 + 4.5y-6 = -10
1.2x + 9.8 -7.5y = 2
2.4x -10 + 4.5y = -10
2.4x x -7.5 -10x-7.5 +4.5y-7.5 = 10x-7.5
-18x +75 -33.75y = -75
1.2x-18x +9.8+75 - 7.5y-33.75y = -75+2
16.8x + 84.8 - 41.25y = 77

Answer by stanbon(75887)

1.2(x+4)-2.5(3y-2) = 2
Multiply thru by 10 to get rid of the decimals:
12(x+4) - 25(3y-2) = 20
12x+48-75y+50 = 20
12x-75y= -78
Divide thru by 3 to get:
4x - 25y = -26
---------------------
(4/5)(3x-5)+ (3/2)(3y-4) = -10
Multiply thru by 10 to get rid of the denominators:
8(3x-5) + 15(3y-4) = -100
24x-40+45y-60 = -100
24x+45y = 0
8x+15y = 0
-----------------------
Solve this system of equations:
1st: 4x - 25y = -26
2nd: 8x + 15y = 0
----------------
Multiply 1st by "2" to get:
3rd: 8x - 50y = -52
------------------
Subtract 2nd from 3rd to get:
-65y = -52
y = (4/5) or 0.8
----------------
Substitute into 2nd to solve for "x"
8x + 15(4/5) =0
8x + 12 = 0
x = -1.5
--------------
Final solution:
x= -1.5
y= 0.8
=========================
Check the answer in 1.2(x+4)-2.5(3y-2) = 2
1.2((-1.5)+4)-2.5(3*0.8-2) = 2
1.2(2.5) - 2.5(0.4) = 2
3 - 1 = 2
2 = 2
===============
Cheers,
Stan H.